IN AN EXCEL FILE USING EXCEL FUNCTIONS CALCULATIONS SHOW FORMULAS IN CELL PLEASE The annual number of industrial accidents occurring in a manufacturing plant is known to follow a Poisson distribution with mean 12. a. What is the probability of observing of observing exactly 12 accidents during the coming year? b. What is the probability of observing no more than 12 accidents during the coming year? c. What is the probability of observing at least 15 accidents during the coming year? d. What is the probability of observing between 10 and 15 accidents (including 10 and 15) during 1) the coming year? Find the smallest integer k such that we can be at least 99% sure that the annual number of accidents occurring will be less than k. e.
Answer #1

Let X ~ Poi( lambda = 12 )

a) Here we want to find P( X = 12)

Let's use excel:

P(X = 12) = "=POISSON(12,12,0)" = 0.1144

b) Here we want to find P( X <= 12) = "=POISSON(12,12,1)" = 0.5760

c) Here we want to find P( X >= 15) = 1 - P( X <= 14) ...( 1 )

Let P(X <= 14) = "=POISSON(14,12,1)" = 0.7720

Plug this value in equation ( 1 ), we get.

P(X >= 15) = 1 - 0.7720 = 0.2280

d) Here we want to find P( 10 <= X <= 15 ) = P(X <= 15) - P(X <= 9) ......( 2 )

P(X <= 15 ) = "=POISSON(15,12,1)" = 0.8444

P(X <= 9 ) = "=POISSON(9,12,1)" = 0.2424

Plug these values in equation ( 1 ), we get.

P( 10 <= X <= 15 ) = 0.8444 - 0.2424 = 0.6020

e) Here we want to find k such that P(X < k) >= 0.99

Let's make cumulative Poisson probability table in excel:

P(x) Е(х) 0 6.14421E-06 6.14421E-06 1 7.37305Е-05 7.98748Е-05 2 0.000442383 0.000522258 3 0.001769533 0.002291791 4 0.005308599 0.007600391 5 0.012740639 0.020341029 6 0.025481277 0.045822307 7 0.04368219 0.089504497 8 0.065523285 0.155027782 9 0.08736438 0.242392162 10 0.104837256 0.347229418 11 0.114367916 0.461597333 12 0.114367916 0.575965249 13 0.105570384 0.681535632 14 0.0904889 0.772024532 15 0.07239112 0.844415652 16 0.05429334 0.898708993 17 0.038324711 0.937033703 18 0.025549807 0.96258351 19 0.01613672 0.978720231 20 0.009682032 0.988402263 5 10 12 13 14 15 16 17 18 19 20 21 23 24 21 0.00553259 0.993934853 We get F(x) >0.99 here. So that the value of k 21 22 0.003017776 0.996952629 ▶▶1 Sheet1 Sheet2 Sheet3

The formulae used on the above excel-sheet are as follows:

P(x) POISSON(A2,12,0) POISSON(A3,12,0) POISSON(A4,12,0) POISSON(A5,12,0) POISSON(A6,12,0) POISSON(A7,12,0) POISSON(A8,12,0) POISSON(A9,12,0) POISSON(A10,12,0) POISSON(A11,12,0) POISSON(A12,12,0) POISSON(A13,12,0) POISSON(A14,12,0) POISSON(A15,12,0) POISSON(A16,12,0) POISSON(A17,12,0) POISSON(A18,12,0) POISSON(A19,12,0) POISSON(A20,12,0) POISSON(A21,12,0) POISSON(A22,12,0) POISSON(A23,12,0) POISSON(A24,12,0) F(x) -B2 -B3+C2 -B4+C3 B5+C -B6+C5 -B7+C6 -B8+C7 4 2 6 4 9 7 10 8 11 9 12 10 13 11 14 12 15 13 16 14 17 15 18 16 19 17 20 18 21 19 22 20 23 21 24 22 B10+C9 -B11+C10 -B12+C11 -B13+C12 B14+C13 -B15+C14 B16+C15 -B17+C16 -B18+C17 B19+C18 B20+C19 B21+C20 -B22+C21 -B23+C22 B24+C23 We get F(x) >- 0.99 h

Therefore answer of part e) is k = 21